# A comprehensive analysis toward the Fermi-LAT source 4FGL J1846.90227 Jets of a proto-planetary nebula producing -rays?

### v1

@@ Abstract
Most of the gamma-ray sources in the Fermi-LAT 14-year Source Catalog are associated with pulsars and blazars. However, unveiling the nature of the still unassociated gamma-ray sources is important for the understanding of high energy emission mechanisms in astrophysical objects. This work presents a comprehensive study toward the region covered by the Fermi source 4FGL J1846.90227 previously suggested to be a blazar or a massive protostar. Using multiwavelength observations, we analyzed several astrophysical objects in the region as possible counterparts of the Fermi-LAT source. Having discarded most of them after a comprehensive analysis, we suggest that the most likely candidate to be such a counterpart is IRAS 184430231. We discovered that this source, previously cataloged as a planetary nebula candidate, actually is a proto-planetary nebula. The radio continuum image at 3 GHz associated with such a nebula allowed us to identify a jet-like structure. Additionally, we identified an associated red-shifted CO molecular outflow and a dense molecular clump in which the source is embedded. We obtained a radio spectral index of for the source, indicating syncrothron emission due to accelerated particles. Thus, we suggest that processes such as proton-proton collisions and relativistic Bremsstrahlung are likely to occur. IRAS 184430231, lying almost at the center of the Fermi confidence ellipse and related to the hard X-ray source 4XMM J184700.4022752, would be the first association between a proto-planeatry nebula and gamma emission.
1 Introduction
@@
Fermi-LAT unassociated sources represent some of the most enigmatic gamma-ray sources in the sky. In the latest data release from the Fermi Gamma-ray Space Telescope (the 4th Fermi LAT 14 yr Catalog, or 4FGL-DR4), containing 7194 gamma-ray sources (Ballet etĀ al., 2023), more than 50% of the suspected Galactic sources (i.e., ) are yet to be identified (Rangelov etĀ al., 2024).
@@
The nature of the gamma-ray sources may be varied, they can be pulsar-wind nebulae (PWNe) and supernova remnants (SNRs) (e.g. Acero etĀ al. 2013 and Cao etĀ al. 2023), active galactic nuclei (AGNs; e.g. Madejski & Sikora 2016), massive young stellar objects (MYSOs; e.g. de OƱa Wilhelmi etĀ al. 2023), X-ray binaries (XRBs; e.g. Kretschmar etĀ al. 2019 and Bahramian & Degenaar 2023), and classic- and symbiotic-novae (Franckowiak etĀ al., 2018). According to the 4FGL catalogs, many of the GeV-emitting PWNe and SNRs are found to be extended LAT sources, many AGNs are found to be variable sources, classical and symbiotic novae are only found to be GeV emitters during outbursts, and MYSOs were argued to be GeV emitters based mainly on positional coincidences.
@@
In short, any source capable of accelerating particles should be considered as a possible responsible of the gamma-ray emission. Thus, given the great variety of objects that can emit at high energies and given the typical spatial extent of the Fermi sources, revealing the nature of unidentified gamma-ray sources requires comprehensive multiwavelength studies both of the possible objects and their associated emission mechanisms.
@@
4FGL J1846.90227 gamma-ray source (first cataloged as 1FGL J1846.8-0233c) was proposed to be associated with the MYSO candidate MSX6C G030.198100.1691 (hereafter MSX G30; Munar-Adrover etĀ al., 2011), based on the spatial correlation between both sources and considering the absence, according to the knowledge at that time, of other possible counterparts. Massive protostars have associated energetic jets and molecular outflows that can produce strong shocks during the interaction with the ambient gas. At these shocks, particles can be accelerated up to relativistic energies producing gamma-ray emission, as some theoretical models predict (e.g., Araudo & RodrĆguez, 2012; Bosch-Ramon etĀ al., 2010).
@@
A decade later, Kerby etĀ al. (2021) analyzed a sample of 174 unknown Fermi sources with a single X-ray/UV/optical counterpart. The authors developed a method to discern between pulsars and blazars, given that the most of the sources of 4FGL catalog are these types of objects. In particular, the authors found two Swift-XRT X-ray sources as possible counterpart of the gamma-ray source 4FGL J1846.90227. The first one, SwXF4 J184650.7022904, was rejected as possible pulsar or blazar because of its extreme photon index in the X-ray band (). Moreover, the authors warned about the positional coincidence (shifted about 4ā³) of this Swift-XRT source with a dim star, suggesting that the high āblazar probability valueā obtained should not be trusted, given that UV/optical emission would not be related to the blazar. On the other hand, the second Swift source identified towards the region, SwXF4 J184651.6022507, whose X-ray flux is three orders of magnitude smaller than the first one, was characterized as a likely blazar.
@@
It is worth mentioning that despite all the studies carried out to date, the exact origin of the -ray emission from 4FGL J1846.90227 remains unknown. In fact, in the last two incremental versions (4FGL-DR3 and 4FGL-DR4) of the fourth Fermi Large Area Telescope (LAT) catalog of -ray sources (Abdollahi etĀ al. 2022 and Ballet etĀ al. 2023, respectively), this source still appears as āunknownā. However, it is important to mention that both catalogs refer to a possible association with the radio source TXS 1844025 mentioned in Honma etĀ al. (2000). Using the Japanese VLBI Network, the authors carried out radio continuum observations at 22 GHz toward a sample of 267 radio sources selected from existing radio surveys. In particular, they did not detect emission at 22 GHz towards the source TXS 1844025. This radio source is in positional coincidence with the infrared source IRAS 184430231, which is cataloged as a planetary nebula (PN) candidate (Yang etĀ al. 2016; Irabor etĀ al. 2018, and references therein).
@@
Summarizing, since the nature of 4FGL J1846.90227 is still unknown, we present a comprehensive and multi-wavelength analysis of all candidate sources in the region to identify definitely the origin of the -ray emission. It is important to highlight that this type of in-depth studies of specific sources are necessary to complement the statistical studies.
@@ 2.1 Radio continuum data
The continuum at 20ācm (1.4 GHz) was extracted from the Multi-Array Galactic Plane Imaging Survey (MAGPIS; Helfand etĀ al., 2006), which has an angular resolution and sensitivity of about 5ā²ā²and 2Ā mJyĀ beam-1, respectively. Even with data from the D configuration, the resulting maps suffer missing flux from large-scale structure (above 1ā²) to which the VLA is insensitive. To correct for this deficiency, the VLA images were combined with images from a 1400 MHz survey made with the Effelsberg 100Ā m telescope (Reich etĀ al., 1990, beam 94).
@@
The radio continuum at 5 GHz was extracted from the Co-Ordinated Radio āNā Infrared Survey for High-mass star formation (CORNISH; Purcell etĀ al., 2008). The survey was conducted using the Karl Jansky Very Large Array in B-configuration, yielding a 15 resolution map with a noise level 0.4 mJy/beam-1. The largest angular scale is 15ā³.
@@ 2.2 Mid-infrared and submillimeter data
The 12CO J=3ā2 data from the James Clerk Maxwell Telescope (JCMT) was extracted from the Canadian Astronomy Data Centre. The PI and proposal ID are Lumsden, S. and M07AU08, respectively. The angular and velocity resolutions are 15 and 0.4 kms-1, respectively. The rms noise level is about 0.1 K.
@@ 2.3 ALMA data
The data cube from the project 2015.1.01312 (PI: Fuller, G.; Band 6) were obtained from the ALMA Science Archive333http://almascience.eso.org/aq/. The single pointing observation for the target was carried out using the telescope configurations with L5BL/L80BL(m) 42/222 in the array. The frequency range used in our analysis goes from 224.3 to 226.3Ā GHz. The angular resolution is 07. The maximum recoverable scale is about 62. The field of view is 25. The frequency and velocity resolutions are 1.1Ā MHz and 1.4Ā km s-1, respectively. The continuum and line (10Ā km s-1Ā averaged) sensitivity are 0.2 and 1.5Ā mJyĀ beam-1, respectively.
Although the data of the project passed the QA2 quality level, which assures a reliable calibration for āscience readyā data, we reprocessed the raw data using CASA 4.5.1 and 4.7.2 versions and the calibration pipeline scripts to check the final images. Particular care was taken with the different parameters of clean the task. The images and spectra obtained from our data reprocessing āafter several runs of the clean task, varying some of its parametersā were very similar to those obtained from the archival data. Briefly, the quality of the images automatically processed by the ALMA team could not be improved.
The task imcontsub in CASA was used to subtract the continuum from the spectral lines using a first-order polynomial. To to that frequency ranges without molecular line emission were carefully selected. The continuum map at 240Ā GHz was corrected for primary beam. Several continuum subtraction tests were performed to ensure a reliable 240Ā GHz continuum map, which involved the selection of different free line regions of the spectrum. It is important to note that the beam size of the 240Ā GHz continuum data in the 12Ā m array provides a spatial resolution of about 0.02Ā pc ( 4000 au) at 5.4Ā kpc, which is the kinematic distance assigned to the MYSO candidate MSX G030.
@@ 2.4 XMM-Newton data
The positional error ellipse of 4FGL J1846.90227 falls within the field-of-view of the XMM-Newton observations 0651680301 and 0823990801, hereafter Obs1 and Obs2, respectively. The observations were pointed to the supergiant fast-X-ray transient IGR J184620223, and the centroid of the Fermi source is located off-set from the center of the X-ray cameras. The duration of Obs1 and Obs2 are 33.4 and 35.7Ā ks, respectively. We used the Science Analysis System (SAS, version 21.0.0 and HEASoft (version 6.33.2) for data reduction and spectral analysis. We applied the standard procedure to obtain a calibrated event file for each observation using the updated set of calibration files (CCF). We use these event files to produce light curves in the high-energy domain () and filtered out periods of high-count rate associated with background flares. We applied additional filtering to include events with FLAG = 0 and PATTERN 12 and 4 for PN and MOS cameras, respectively. In Table 1 we report a summary of the X-ray observations.
@@
0651680301 | MOS1 | Full Frame | 30.8 |
(Apr. 18, 2011) | MOS2 | Full Frame | 31.6 |
PN | Full Frame | 15.1 | |
0823390801 | MOS1 | Small Window | 31.3 |
(Oct. 22, 2018) | MOS2 | Timing | 31.1 |
PN | Full Frame | 14.5 |
@@ 3 Results
Fermi-LAT sources typically have positional errors of a few arcminutes, which implies that there are often several candidates to be the counterpart of the high energy source. Figureā1 shows a three color image with 8 and 24 m, and 20 cm emissions represented in green, red, and blue, respectively. The center and the 95% confidence ellipse of the Fermi source 4FGL J1846.90227 are represented by the yellow cross and the dashed-yellow ellipse, respectively. We carried out the search for the Fermi counterpart in the region within such an ellipse.
Figureā2-left panel shows the whole field of view of the XMM-Newton observations with the X-ray energy band between 0.5 keV and 8.0 keV. This image was obtained by combining the 6 images and exposure maps o the mos1, mos2 and pn cameras of each observation using the emosaic task of SAS. We applied a Gaussian smoothing to increase the S/N ratio. Figureā2-right panel shows a two-color image of the X-ray soft/medium and hard emissions between 0.5 and 2.0 keV (red) and 2.0 and 8.0 keV (green), respectively. No hint of diffuse emission coincident with the Fermi source and some point sources can be clearly observed. The brightest X-ray point source, by far, coincident with the Fermi positional error ellipse is 4XMM J184650.6022907 (associated with SwXF4 J184650.7022904), a soft source with no counts detected above 3.0 keV. Additionally are indicated the soft X-ray source 4XMM J184651.2022504 (associated with SwXF4 J184651.6022507) and the hard X-ray source 4XMM J184700.4022752 (associated with the PN candidate IRAS 184430231). These sources are analyzed below.
The region considered for the analysis includes four centimeter radio continuum sources (see Fig.ā1 and Fig.ā3-left). G030.21300.156 (hereafter G30.21) is cataloged as an HiiĀ region with a radio recombination line at 100Ā km s-1, which corresponds to the near kinematic distance of about 5.4Ā kpc (Anderson etĀ al., 2015). Southwards G30.21, we identify an uncatalogued extended radio continuum source. It is worth noting that both radio sources, which seem not to be connected, exhibit an evident flattening facing a conspicuous dust filament seen at 8 and 24 m (see Fig.ā1 and Fig.ā3-left). It is possible that both radio sources compress such a dust filament generating its flattened morphology. It is important to mention that Anderson etĀ al. (2015) used Green Bank Telescope radio observations, whose angular resolution of about 87 prevented them from resolving both radio sources, which were considered as a single one.
@@
On the other hand, it can also be appreciated two additional compact radio sources in the region, IRAS 184430231 and MWP1G030218001825, which have been cataloged as a possible planetary nebula, and a compact HiiĀ region (Anderson etĀ al., 2011), respectively. Given that the nature of IRAS 184430231 is still an open issue and it is associated with the X-ray source 4XMM J184700.4022752, it deserves more analysis in the context of this work (see Sect.ā3.5). In such a context, the three X-ray sources indicated in Fig.ā2-right panel are analyzed in detail.
In addition, the possible contribution to the -ray emission of the MYSO candidate MSX G30 (see Fig. 3), which is embedded in the dust clump AGAL G030.19780.1683, is analyzed (see Sect.ā3.2).

### v2

# A comprehensive analysis towards the Fermi-LAT source 4FGL J1846.90227: jets of a proto-planetary nebula producing -rays?

@@ Abstract
Most of the -ray sources in the Fermi-LAT 14-year Source Catalogue are associated with pulsars and blazars. However, unveiling the nature of the still unassociated -ray sources is important for the understanding of high energy emission mechanisms in astrophysical objects. This work presents a comprehensive study towards the region covered by the Fermi source 4FGL J1846.90227, previously suggested to be a blazar and a massive protostar. Using multiwavelength observations, we analysed several astrophysical objects in the region as possible counterparts of the Fermi-LAT source. Having discarded most of them after a detailed and comprehensive analysis, we suggest that the most likely candidate to be such a counterpart is IRAS 184430231, a likely proto-planetary nebula. We found that the radio continuum emission at 3 GHz of IRAS 184430231 shows a compact source related to faint emission with jet-like morphology. Additionally, we identified an associated red-shifted CO molecular outflow. Using data from several catalogues, we obtained radio spectral index values ranging from 0.57 to 0.39 for IRAS 184430231, indicating syncrothron emission due to particles accelerated by the jets. We point out that these jets could explain the -ray emission through mechanisms such as proton-proton collisions and relativistic Bremsstrahlung. IRAS 184430231, lying almost at the centre of the Fermi confidence ellipse and related to the hard X-ray source 4XMM J184700.4022752, would be the first association between a proto-planetary nebula and -ray emission.
keywords:
Gamma rays: ISM ā (Stars:) binaries: symbiotic ā (ISM:) planetary nebulae: general ā ISM: jets and outflows1 Introduction
@@
Fermi-LAT unassociated sources represent some of the most enigmatic -ray sources in the sky. In the latest data release of the Fermi Gamma-ray Space Telescope (the 4th Fermi LAT 14 yr Catalogue, or 4FGL-DR4), containing 7194 -ray sources (Ballet etĀ al., 2023), more than 50% of the suspected Galactic sources (i.e., ) are yet to be identified (Rangelov etĀ al., 2024).
@@
The nature of the -ray sources may be varied, they can be pulsar-wind nebulae (PWNe) and supernova remnants (SNRs) (e.g. Acero etĀ al. 2013 and Cao etĀ al. 2023), active galactic nuclei (AGNs; e.g. Madejski & Sikora 2016), massive young stellar objects (MYSOs; e.g. de OƱa Wilhelmi etĀ al. 2023), X-ray binaries (XRBs; e.g. Kretschmar etĀ al. 2019 and Bahramian & Degenaar 2023), and classic- and symbiotic-novae (Franckowiak etĀ al., 2018). According to the 4FGL catalogues, many of the GeV-emitting PWNe and SNRs are found to be extended LAT sources, many AGNs are found to be variable sources, classical and symbiotic novae are only found to be GeV emitters during outbursts, and MYSOs were argued to be GeV emitters based mainly on positional coincidences.
@@
In short, any source capable of accelerating particles should be considered as a possible responsible of the -ray emission. Thus, given the great variety of objects that can emit at high energies and given the typical spatial extent and positional error of the Fermi sources, revealing the nature of unidentified -ray sources requires comprehensive multiwavelength studies both of the possible objects and their associated emission mechanisms.
@@
4FGL J1846.90227 is a -ray source (first catalogued as 1FGL J1846.8-0233c) with no spatial extent detected, then classified as a point source. It was proposed to be associated with the MYSO candidate MSX6C G030.198100.1691 (hereafter MSX G30; Munar-Adrover etĀ al., 2011), based on the spatial correlation between both sources and considering the absence, according to the knowledge at that time, of other possible counterparts. Massive protostars have associated energetic jets and molecular outflows that can produce strong shocks during the interaction with the ambient gas. At these shocks, particles can be accelerated up to relativistic energies producing -ray emission, as some theoretical models predict (e.g., Araudo & RodrĆguez, 2012; Bosch-Ramon etĀ al., 2010).
@@
A decade later, Kerby etĀ al. (2021) analysed a sample of 174 unknown Fermi sources with a single X-ray/UV/optical counterpart. The authors developed a method to discern between pulsars and blazars, given that the most of the sources of 4FGL catalogue are these types of objects. In particular, the authors found two Swift-XRT X-ray sources as possible counterpart of the -ray source 4FGL J1846.90227. The first one, SwXF4 J184650.7022904, was rejected as possible pulsar or blazar because of its extreme photon index in the X-ray band (). Moreover, the authors warned about the positional coincidence (shifted about 4ā³) of this Swift-XRT source with a dim star, suggesting that the high āblazar probability valueā obtained should not be trusted, given that UV/optical emission would not be related to the blazar. On the other hand, the second Swift source identified towards the region, SwXF4 J184651.6022507, whose X-ray flux is three orders of magnitude smaller than the first one, was characterised as a likely blazar.
@@
It is worth mentioning that despite all the studies carried out to date, the exact origin of the -ray emission from 4FGL J1846.90227 remains unknown. In fact, in the last two incremental versions (4FGL-DR3 and 4FGL-DR4) of the fourth Fermi Large Area Telescope (LAT) catalogue of -ray sources (Abdollahi etĀ al. 2022 and Ballet etĀ al. 2023, respectively), this source still appears as āunknownā. However, it is important to mention that both catalogues refer to a possible association with the radio source TXS 1844025 mentioned in Honma etĀ al. (2000). Using the Japanese VLBI Network, the authors carried out radio continuum observations at 22 GHz towards a sample of 267 radio sources selected from existing radio surveys. In particular, they did not detect emission at 22 GHz towards the source TXS 1844025. This radio source is in positional coincidence with the infrared source IRAS 184430231, which is catalogued as a planetary nebula (PN) candidate (Yang etĀ al. 2016; Irabor etĀ al. 2018, and references therein).
@@
Summarising, since the nature of 4FGL J1846.90227 is still unknown, we present a comprehensive and multi-wavelength analysis of all candidate sources in the region to identify definitely the origin of the -ray emission. It is important to highlight that this type of in-depth studies of specific sources are necessary to complement the statistical studies.
@@ 2.1 Radio continuum data
The continuum at 20ācm (1.4 GHz) was extracted from the Multi-Array Galactic Plane Imaging Survey (MAGPIS; Helfand etĀ al., 2006a), which has an angular resolution and sensitivity of about 5ā²ā²and 2Ā mJyābeam-1, respectively. Even with data from the D configuration, the resulting maps suffer missing flux from large-scale structure (above 1ā²) to which the VLA is insensitive. To correct for this deficiency, the VLA images were combined with images from a 1400 MHz survey made with the Effelsberg 100Ā m telescope (Reich etĀ al., 1990, beam 94).
@@
The radio continuum at 5 GHz was extracted from the Co-Ordinated Radio āNā Infrared Survey for High-mass star formation (CORNISH; Purcell etĀ al., 2008). The survey was conducted using the Karl Jansky Very Large Array in B-configuration, yielding a 15 resolution map with a noise level 0.4 mJyābeam-1. The largest angular scale is 15ā³.
@@ 2.2 Mid-infrared and submillimeter data
The 12CO J=3ā2 data from the James Clerk Maxwell Telescope (JCMT) was extracted from the Canadian Astronomy Data Centre. The PI and proposal ID are Lumsden, S. and M07AU08, respectively. The angular and velocity resolutions are 15 and 0.4 kmās-1, respectively. The rms noise level is about 0.1 K.
@@ 2.3 ALMA data
The data cube from the project 2015.1.01312 (PI: Fuller, G.; Band 6) were obtained from the ALMA Science Archive333http://almascience.eso.org/aq/. The single pointing observation for the target was carried out using the telescope configurations with L5BL/L80BL(m) 42/222 in the array. The frequency range used in our analysis goes from 224.3 to 226.3Ā GHz. The angular resolution is 07. The maximum recoverable scale is about 62. The field of view is 25. The frequency and velocity resolutions are 1.1Ā MHz and 1.4Ā kmās-1, respectively. The continuum and line (10Ā kmās-1Ā averaged) sensitivity are 0.2 and 1.5Ā mJyĀ beam-1, respectively.
Although the data of the project passed the QA2 quality level, which assures a reliable calibration for āscience readyā data, we reprocessed the raw data using CASA 4.5.1 and 4.7.2 versions and the calibration pipeline scripts to check the final images. Particular care was taken with the different parameters of clean task. The images and spectra obtained from our data reprocessing āafter several runs of the clean task, varying some of its parametersā were very similar to those obtained from the archival data. Briefly, the quality of the images automatically processed by the ALMA team could not be improved.
The task imcontsub in CASA was used to subtract the continuum from the spectral lines using a first-order polynomial. To that frequency ranges without molecular line emission were carefully selected. The continuum map at 240Ā GHz was corrected for primary beam. Several continuum subtraction tests were performed to ensure a reliable 240Ā GHz continuum map, which involved the selection of different free line regions of the spectrum. It is important to note that the beam size of the 240Ā GHz continuum data in the 12Ā m array provides a spatial resolution of about 0.02 pc ( 4000 au) at 5.4 kpc, which is the kinematic distance assigned to the MYSO candidate MSX G030.
@@ 2.4 XMM-Newton data
The positional error ellipse of 4FGL J1846.90227 falls within the field-of-view of the XMM-Newton observations 0651680301 and 0823990801, hereafter Obs1 and Obs2, respectively. The observations were pointed to the supergiant fast-X-ray transient IGR J184620223, and the centroid of the Fermi source is located off-set from the centre of the X-ray cameras. The duration of Obs1 and Obs2 are 33.4 and 35.7Ā ks, respectively. We used the Science Analysis System (SAS, version 21.0.0) and HEASoft (version 6.33.2) for data reduction and spectral analysis. We applied the standard procedure to obtain a calibrated event file for each camera of each observation using the updated set of calibration files (CCF). We use these event files to produce light curves in the high-energy domain () and filtered out periods of high-count rate associated with background flares. We applied additional filtering to include events with FLAG = 0 and PATTERN 12 and 4 for EPIC-pn and EPIC-mos cameras, respectively. In Table 1 we report a summary of the X-ray observations.
@@
0651680301 | EPIC-mos1 | Full Frame | 30.8 |
(Apr. 18, 2011) | EPIC-mos2 | Full Frame | 31.6 |
EPIC-pn | Full Frame | 15.1 | |
0823390801 | EPIC-mos1 | Small Window | 31.3 |
(Oct. 22, 2018) | EPIC-mos2 | Timing | 31.1 |
EPIC-pn | Full Frame | 14.5 |
@@ 3 Results
Fermi-LAT sources typically have positional errors of a few arcminutes, which implies that there are often several candidates to be the counterpart of the high-energy source. Figureā1 shows a three colour image with 8 and 24 m, and 20 cm emissions represented in green, red, and blue, respectively. The centre and the 95% positional error confidence ellipse of the Fermi source 4FGL J1846.90227 are represented by the yellow cross and the dashed-yellow ellipse, respectively. We carried out the search for the Fermi counterpart in the region within such an ellipse.
Figureā2-left panel shows the whole field of view of the XMM-Newton observations in the X-ray energy band between 0.5 keV and 8.0 keV. This image was obtained by combining the 6 images and exposure maps of the EPIC-mos1, EPIC-mos2 and EPIC-pn cameras of each observation using the emosaic task of SAS. We applied a Gaussian smoothing to increase the S/N ratio. Figureā2-right panel shows a two-colour image of the X-ray soft/medium and hard emissions between 0.5 and 2.0 keV (red) and 2.0 and 8.0 keV (green), respectively. No hint of diffuse emission coincident with the Fermi source and some point sources can be clearly observed. The brightest X-ray point source, by far, coincident with the Fermi positional error ellipse is 4XMM J184650.6022907 (associated with SwXF4 J184650.7022904), a soft source with no counts detected above 3.0 keV. Additionally are indicated the soft X-ray source 4XMM J184651.2022504 (associated with SwXF4 J184651.6022507) and the hard X-ray source 4XMM J184700.4022752 (associated with the PN candidate IRAS 184430231). These sources are analysed below.
The region considered for the analysis includes four centimeter radio continuum sources (see Fig.ā1 and Fig.ā3-left). G030.21300.156 (hereafter G30.21) is catalogued as an HiiĀ region with a radio recombination line at 100Ā kmās-1, which corresponds to the near kinematic distance of about 5.4Ā kpc (Anderson etĀ al., 2015). Southwards G30.21, we identify an uncatalogued extended radio continuum source. It is worth noting that both radio sources, which seem not to be connected, exhibit an evident flattening facing a conspicuous dust filament seen at 8 and 24 m (see Fig.ā1 and Fig.ā3-left). It is possible that both radio sources compress such a dust filament generating its flattened morphology. It is important to mention that Anderson etĀ al. (2015) used Green Bank Telescope radio observations, whose angular resolution of about 87 prevented them from resolving both radio sources, which were considered as a single one.
@@
On the other hand, it can also be appreciated two additional compact radio sources in the region, IRAS 184430231 and MWP1G030218001825, which have been catalogued as a possible planetary nebula, and a compact HiiĀ region (Anderson etĀ al., 2011), respectively. Given that the nature of IRAS 184430231 is still an open issue and it is associated with the X-ray source 4XMM J184700.4022752, it deserves more analysis in the context of this work (see Sect.ā3.5). In such a context, the three X-ray sources indicated in Fig.ā2-right panel are analysed in detail.
In addition, the possible contribution to the -ray emission of the MYSO candidate MSX G30 (see Fig. 3), which is embedded in the dust clump AGAL G030.19780.1683, is analysed (see Sect.ā3.2).